OpenStudy (anonymous):

Use the substitution method to solve the system of equations. Enter your answer as an ordered pair. y = 9x + 4 x = 6

OpenStudy (anonymous):

x = 6. substitute that in the other equation. and find the value of y.

OpenStudy (anonymous):

The problem should look like this when a value is known. y = 9(6)+4 complete the problem to solve for y

OpenStudy (anonymous):

(6,58)

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

y = 9*6+4 y = 54+4 y = 58